OpenStudy (anonymous):

Ahmed is taking orders for lunch. A slice of pizza cost $2.1 and a chicken sandwich costs $3.1. He collects $22.8 for the group of 8 people. Let X = number of pizza orders. Let Y = number of chicken orders. Assuming each person orders only one item, How many people ordered a chicken sandwich?

OpenStudy (freckles):

well we are given the number of people is 8 and the people have been divided into groups X and Y so we know X+Y=8 since everyone ordered one thingy be it chicken or pizza we are also given that pizza costs 2.1 and chicken costs 22.8 so the total the group X spends will be 2.1*X and total the group Y spend will be 3.1*Y and we know the totals of those totals should be equal to 22.8 so we know 2.1*X+3.1*Y=22.8

OpenStudy (freckles):

You have a system of two linear equations

OpenStudy (freckles):

You could chose to use substitution or elimination to solve.

OpenStudy (anonymous):

still dont get it sorry

OpenStudy (freckles):

I'm sure you solved a system of equations before since this is a word problem...

OpenStudy (freckles):

X+Y=8 <---solve this for X replace the X in 2.1*X+3.1*Y=22.8 with what you found X to be in the first eqution

OpenStudy (anonymous):

got it